OpenStudy (anonymous):

prove the following identity: cos 2x+cos x+1/sin 2x+sin x=cot x thanks

OpenStudy (anonymous):

\[cos2x+cos x+1/sin2x+sin x=cot x\]

OpenStudy (anonymous):

Simplifying the function and using trigonometric identities, you can prove the identity.

OpenStudy (mathstudent55):

Tis is what you wrote: \( \cos2x+\cos x+\dfrac{1}{\sin2x}+\sin x=\cot x\) You can't prove that identity because it isn't one. Did you mean this? \( \dfrac{\cos2x+\cos x+1}{\sin2x+\sin x} = \cot x \)

OpenStudy (anonymous):

yes it is

OpenStudy (mathstudent55):

Here are two identities that may be useful. \(\cos 2x = \cos^2 x - sin^2 x\) \(\sin 2x = 2 \sin x \cos x\)

OpenStudy (anonymous):

\[\frac{ \cos 2x+1+\cos x }{ \sin 2x+\sin x }=\frac{ 2 \cos ^2x+\cos x }{ 2 \sin x \cos x+\sin x }=\frac{ \cos x \left( 2 \cos x+1 \right) }{ \sin x \left( 2 \cos x+1 \right) } =?\]